OpenStudy (anonymous):

The bicycle store has moutain bikes and racing bikes on the showroom floor at a ratio of 3 to 5. If there are a total of 72 bicycles on the showroom floor, how can you find the number of racing bikes? Use your method to find the answer.

OpenStudy (anonymous):

Let x be the number of bikes. In this case 3x + 5x = 72. 8x = 72 so x = 9. Because the ratio of mountain to cycling is 3 to 5. You would times 5 * 9 = 45. So in this store there are 45 racing bikes and 72 - 45 = 27 mountain bikes.

OpenStudy (anonymous):

@_beautygee are you there?

OpenStudy (anonymous):

@Boblovesmath yea thnx!